Session: Nov 1998 (session 4)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Sergio Mariotti at Mon Nov 16 09:43:23 MET 1998) Success
Everything appear OK

Onsala (A. Polatidis at Mon Nov 16 10:29:27 MET 1998) Success
No known problems

Torun (K.Borkowski at Mon Nov 16 14:20:07 MET 1998) Minor failures
Problems with stability of the receiver started at about 19:50 UT. Telescope then stowed at 20:17 and observations resumed at 22:20.

Effelsberg (dgraham at Mon Nov 16 16:47:37 MET 1998) Minor failures
no problems noticed yet

Jodrell (Alastair Gunn at Wed Nov 18 11:28:41 MET 1998) Success
No known problems.

Calibration (Jean\-Francois Desmurs at Tue Nov 24 19:08:04 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov98 directory under the name: gb028.antab.Read the comments in the file to get further information.